Sunday 7/7/02

Dave, Steve, Jim and Andrew arrive in Chile. Steve decides he doesn'ít want to go home and throws his airline tickets away. They collect our Toyota pickup and drive up to the observatory. A very cheesy radio station is discovered - there ain't much choice up in the Andes. The team arrive on the mountain and inspect the CIRPASS shipment. Everything looks to be in good condition. A huge condor says hello near the 4m at CTIO. Andrew eats his 17th meal since leaving Cambridge at CTIO.

Monday 8/7/02

Work starts on unpacking and assembling CIRPASS. The dewar vacuum vessel is put together without optics etc. to check vacuum integrity. It works fine. Dave starts assembling and aligning the camera optics. The freezer enclosure starts to come together. Some mods are needed to the castors. The detector is successfully checked out while its warm. Jim has an impromptu video-con with Doug Simons in Hawaii

Tuesday 9/7/02

Ian arrives. Work continues on assembling the freezer. The detector focus mechanism is tested out and appears to be working as it was in Cambridge. The cheesy radio station is christened "Radio 2 and a half". If you ever wondered what happened to stuff once it became too cheesy for Radio Two now you know.

Wednesday 10/7/02

Dave locks himself out of his room. Jim takes pity on him and lends him some clothes, sparing him embarrassment at breakfast. Ian overrules Steve's "no cleaning teeth after breakfast" policy. FPU is unpacked and a start is made to prepare it for installation on the telescope. The castor mods are completed and the freezer is wheeled to its allocated position for tests with the Celestron. The compressor and evaporator are mated together. The glycol hoses are made up and installed. Steve subjects us to the delights of X-ray specs and Polly Styrene. The camera optics are installed in the dewar. The filter wheel is successfully checked out.

Thursday 11/7/02

The heat shields are installed and some better spec O-rings fitted. The fully assembled dewar is put on the vacuum pump. The Celestron with its Meade guider and heavy duty tripod mount is unpacked and assembled. The local refrigeration contractors arrive and start work on getting our freezer working. They check out all the plumbing and fill the system with refrigerant. The freezer enclosure is put together empty to prepare for a cool-down test. We all go back to Cerro Tololo for dinner then Andrew, Dave and Ian go back to Cerro Pachon to test out the Celestron. We find that the accessible sky is +4

Friday 12/7/02

The FPU is taken up to the observing floor and test fitted to the balast weight assembly (BWA). With a bit of adjusting and rearranging we make it fit. Work continues on the freezer. The cold test is successful and we reach -40oC as required. This is using the UK plant operating at 380V between phases and 50Hz. It had been working at 430V between phases back in Cambridge so a major uncertainty has now been resolved.

Saturday 13/7/02

We all head off to Sunny La Serena for a bit of R&R. We get there at about 1100. Sunny La Serena turns out to be cold foggy La Serena. We have a couple of holiday cabins near the sea front but we can't see the sea for the fog. The swimming pool looks nice but unfortunately we've all forgotten to bring our ice skates. We go for a drive round trying to find the town centre which somehow eludes us. However we do find a supermarket and buy a trolley load of beer. We decide to have a light lunch and a beer while we plan the rest of the day.

So after returning to our accomodation we go to the grill next door (deciding not go to the Kamikaze Pub complete with aeroplane sticking out of the roof). We are the only guests but it looks nice and the staff are keen and friendly. While drinking our beers we order huge piles of food. Why we do this is not clear as we are all eating way too much at Cerro Tololo. The waiter brings two fine bottles of Chilean red for us to choose between. We opt for "Cabillo Loco" - the cheaper of the two. It really is an excellent wine. As we eat, more people come in and a bit of atmosphere starts to develop. When our wine bottle is empty the waiter asks enthusiastically if we'd like another one. We decide to ask how much it actually costs. When we find out we order a Cabinet Sauvignon. It's very nice but not quite in the same league as the crazy horse. All in all the meal was excellent and we pay and leave having had a thoroughly good time. Even though the crazy horse turns out to be nearly half the total cost for the five of us the overall cost is great value compared to UK prices for such fine food and good service. As we leave the waiter asks when we're coming back.

We decide that we need a driver and a navigator to get to el Centro so we get a cab. We find that we'd been close in our earlier attempt. Once there we go shopping. Ian buys an enormous metal crucifix sculpture. Steve buys a head made of rock. Then we go for a cup of tea where we meet a local chap who is very friendly as indeed are his two young lady friends who join him. Unfortunately, because of our poor spanish we have no idea what they are taking about.

We return to our chilly cabins and watch telly. We vegetate in front of Predator, Dr Strangelove and Silence of the Lambs before going to bed with our clothes on.

Sunday 14/7/02

We are awoken by a cheery lady yelling something in Spanish - our breakfast is ready. We load all our wares into the truck, check out and go off to el Centro again to buy more tourist junk. Steve buys some more rock heads.

On the way back to Tololo we stop at the big dam for a stroll. Away from the coast there is no fog and its nice and sunny. Then we go to Vicuna for lunch. Vicuna is a delightful place with a lovely town square. We have a very nice lunch at a restaurant in the corner of the squeare which is full of locals enjoying their Sunday lunch.

After exploring yet more tourist junk outlets we visit the bug museum. Our 40p entrance fee proves to be a bargain as we see beetles the size of mobile phones and spiders as big as your out-stretched hand.

Monday 15/7/02

The kanya frame for the big spectrograph optics is put into the freezer. The primary mirror is then put in to its operational position. This is always a nerve-wracking procedure but it all goes well. The FPU is removed from the BWA and brought down from the observing floor for tests.

Tuesday 16/7/02

The dewar is filled with LN2 for the first time. Later it is man-handled in to the cold room and the enclosure assembly is completed including the missing walls, roof panels and doors. The IFU slit and the J mask mirror and mask focus mechanism are added.

Wednesday 17/7/02

The rest of the optics are put into the cold room and aligned. The detector checks out fine at LN2 temps. The mechanisms are wired up and tested. We hear that Steve's tickets have been found!

After dinner the complete Celestron assembly is taken outside and the IFU is fitted. Most of the time spent is used to check acquisition and alignment procedures. A few brightish stars are observed. The freezer was left switched off at this stage.

Thursday 18/7/02

It's Andrew's birthday! Unfortunately nobody has bought him a present! Steve receives a present though - his airline tickets.

Further work is done to prepare for switching on the freezer including thermocouples, TV cameras, lights, dewar insulation, window fan etc. The FPU is tested and a fault is identified which is eventully tracked down to a faulty card.

The wheels on the compressor are changed to 4 swivel types in preparation for the move up to the dome floor. We listen to BBC radio 2 on the web and Andrew says he likes it - clearly another birthday has taken its toll. Or perhaps its just an improvement on radio 2.5.

The freezer is switched on at midnight. With a bit of fine tuning it works well. We leave at about 2am with the temperature inside the freezer heading in the right direction.

We have a little celebration of Andrew's birthday back at our accomodation on Tololo. Highlights include some fizzy drinks and Steve's theory of 4-celled embryos.

Friday 19/7/02

The freezer gets to its set temperature of -40 at about 11:15am. The faulty card in the FPU control unit is swapped for the "grating turn" one from the cold room control unit. Grating turn is seldom if ever used and won't be missed.

Saturday 20/7/02

We had planned to go straight from Tololo to La Serena but a power failure up at the telescope at 5am forces us to go back and check the freezer which was reporting a fault condition. Normally the freezer comes back on once power is restored but on this occassion it didn't. It turns out that an overload switch on our plant tripped out when the power failed. The mains probably went up before it went down. Reseting the switch fixes the problem and the freezer is set running again starting from about -12oC.

We arrive in La Serena at about 4pm. The Hotel Francisco de Aguirre where we are staying is very nice. There's a scary moment when our bad spanish nearly results in Steve and Ian sharing a double bed. We go out shopping. Steve needs even more rock heads. We faff about where we should eat. The weather starts to get cold and wet. Steve makes an executive decision and chooses a delightful outdoor restaurant with a view of the taxi rank, booming pop music, holes in the table cloth and a man using a nailgun. We manage to persuade Steve that we should only have a beer.

We go back to the hotel to dump our souvenirs. We phone Helena and arrange to meet her and Marie-Claire at El Muelle on the seafront. We get there by taxi and have a great evening. Andrew is relieved to get some food at last. The weather starts to get worse and by the time we leave El Muelle it's absolutely chucking it down. Steve decides to walk the 3 miles back to the hotel while the rest of us go to a Salsa club.

Sunday 21/7/02

Steve, Jim and Dave are due to fly out today on their long journey home. But to add a bit of tension to the plot our pickup truck has a puncture on the way to the airport. We park in the middle of a dual carriageway to change the wheel. Eventually we find all the things we need to complete the job. While we're mid-frenzy Ian says "er, should we really be blocking the main access to the fire station?". Two policemen wander by and do nothing as if a group of mad British astronomers often repair their punctures in the path of a speeding 20 ton truck on a mission.

We arrive at the airport with loads of time to spare and Andrew and Ian say their farewells to Dave, Steve and Jim. The CIRPASS team has suddenly become the CIRPASS duo. We try calling at the car rental place to get the tyre fixed but it's shut. We drive back up to Tololo giving a ride to a young lad who is walking up the dirt road to the ranch in the mountains. What was rain in La Serena was of course snow at Tololo. We get the puncture repaired at the Tololo garage.

Since leaving the summit yesterday we have rethought our strategy for preparing for our run. The background values in electrons/sec/pix are too high and we consider the possibility that we might have to go inside the dewar to get things fixed. Also, the bad weather has caused some havoc with the phone lines and the power supply and the whole engineering schedule for Gemini. We radio Hernan and ask him to fill the dewar and switch off the freezer. Ram phones at midnight to say his connection at Paris was late so he has to spend an extra day there. There are worse places to be stuck for a day.

Monday 22/7/02

We need a vehicle with snow chains and good tyres to get to Pachon and we don't trust our pickup. We ask if we can get a ride over with someone and we're offered the hire of an AURA pickup complete with CD player. Horray, no more Radio 2.5! We get snow chains fitted and off we go listening to Travis.

At Pachon we find the snow is much deeper and people have worked very hard to clear the road. It's also extremely windy and cold. We have to have one of our snowchains repaired. Everything is well with our equipment and we speed up the warming up process by using the freezer defrost heaters. We test that our light-proofing inside the dewar is working properly by seeing how much light gets to the detector when all the internal freezer lights are switched on. The test indicates the light-proofing is good. We are told we have to evacuate the site at 1530 so we have to stop doing tests and prepare to leave. We get the inside of the freezer to ambient temperature and leave the end door open before heading back to Tololo.

Tuesday 23/7/02

A lot of snow has fallen over night and continues to come down in blizzard conditions. Andrew has a lie-in. We struggle between our house and the other buildings on site by foot while our truck gets buried under more and more snow. We talk with Steve by phone about various extra bits of kit we need out here fairly quickly and he gets them on their way in double time. Nobody goes over to Pachon today.

Wednesday 24/7/02

We awake to the clearest blue skies ever and brilliant white mountain scenes. We are hopeful that we might get to Pachon. However the snow at Tololo is deep so we are not suprised to hear that its very bad at Gemini-S. Ram arrives at lunchtime after a very long trip fom India. The road to Pachon is not cleared until late afternoon and again nobody goes over to the Gemini telescope today. Instead we go over to the Tololo 4-m for a look-see.

Thursday 25/7/02

We manage to get over to Pachon at last. The dewar has warmed up so we start getting the dewar cold again. To improve the background values we need to pump on the LN2 can so we rearrange the plumbing. We have a conference with Manuel, Hernan and Gustavo and decide to move the first night to Thurs 1st instead of wed 31st. This is mainly to give us time should we decide to get inside the dewar. The roads are still very bad so we leave at 1730. First tests of the LN2 pumping plan look good and start to get to grips with the high background problem.

Friday 26/7/02

We start to pump on the LN2 can in earnest now that the dewar has had a chance to cool. The temperature of the nitrogen in the can falls drastically and the background starts to fall too. We take the FPU up to the observing floor and put it into the BWA. The whole assembly is then put on the telescope. The crew decide that the thick steel plate covering the hole where the light comes out should be removed. So we take the BWA off, remove the plate (with a very loud clang) and put the BWA back on again.

At 1730 we switch the freezer on. The stuff Steve shipped arrives into Santiago a day late and gets to La Serena in the evening. So it gets stuck in a warehouse until monday morning. After dinner we monitor the performance of the instrument as it cools. We leave at 2300 optimistic that things are heading in the right direction.

Saturday 27/7/02

Continuing to monitor the cool-down shows that pumping on the LN2 can has worked extremely well and the background values are returning to within spec. We are very relieved by this because now we don't have to open up the dewar. With the instrument almost down to equilibrium temperatures we do some Celestron observing including taking a look at the Lagoon Nebula. We don't finish until 3am in the morning. It's all looking extremely good and we are really looking forward to swapping our $4000 8-inch telescope for a $85 million dollar 8-metre telescope.

Sunday 28/7/02

We make definitive background measurements. For observations in the range 1.00-1.65 microns the background is about 0.15 e/sec/pix. We also do various focus calibrations and check the relative throughputs of the filters. We make sure we've got things ready for the big move to the azimuth platform on monday morning. At 1800 we switch off the freezer. After dinner we start to warm it up with the evaporator heaters.

Monday 29/7/02

Arriving at the telescope we find out that the dome was open during the night while the conditions were very windy and humid. Consequently there's a lot of dust to clean off the telescope structure and mirror by the Gemini staff which delays us moving CIRPASS up to the observing floor. In any case we need a bit of time to prepare including puttong 10mm shims in the castor assemblies to provide a bit more ground clearance. We start to move our computers in to the control room. Rob arrives at Pachon around lunchtime. Anthony arrives at Tololo at lunchtime - wrong mountain but he does somehow manage to get a very nice room in the Dining room Dorm.

After lunch we wheel CIRPASS out of the lab towards the big lifting platform which takes it up to level 4. With one person on each corner it wheels along quite easily and the whole process of getting it on to the azimuth platform is remarkably painless.

Ram gets moved out of house 5 to house 1. Rob is also given a room in house 1. Rob hasn't slept for a very long time and we drop him off outside the house at the first house at the end of our row of houses and point. Ram attempts to move from number 5 to number one. He discovers his key doesn't fit the house at the end and that in fact the house at the end is number two. He eventually finds house number one about 1 km away near the garage. There's no sign of Rob there. We eventually figure out that he's in the wrong house and the wrong bed and we wake him up and drag him over to his proper home. At least he didn't eat the porridge.

Tueday 30/7/02

We finish mounting the FPU on the telescope and get the focal plane in the nominal 300mm backfocus position. Work continue on connecting everything up including the 3 phase supply and the cooling lines for the freezer. Ian complains that we're scattered about all over the place and gets a reputation for being a bit pushy with the locals.

Rachel arrives at Pachon at 1800. We eat at "20 units" and go back to the telescope to watch over the shoulders of the folks using the telescope so we can learn what to do when we get on. We also want to shift ourselves on to a night schedule. The weather's not too good though so we can continue moving our stuff around and setting everthing up ready to observe.

Wednesday 31/7/02

Ram and Rob get moved back to house number 5. We arrive at the telescope at 1300. We make the final preparations. We mount the IFU on the telescope and move the telescope around to check that the fibres dangle in a benign way which of course they do. We make final checks and switch on the freezer and start pumping on the dewar. We get an additional truck to ferry round the Cambridge hoards.

We've been struggling to get our computers in the control room to talk properly to the cortex cards which drive the 15 stepper motor mechanisms. We have to adopt a different system configuration to get it to work but at last we're ready to go.

Officially our first night is tomorrow but as we're ready we're allowed to borrow a bit of time. So tonight we have first light - the first IR IFU on an 8-m telescope. Randy acquires a 4th mag star and puts it in the centre of the telescope field of view. We take an exposure with CIRPASS and it's there on the IFU - in focus and just 1.5 arcsec off centre! We use an hour and a half of time and do some basic things like moving to another star, checking the focus and working out where's north, south, east and west. Weather conditions and seeing are not paticular good.

Thursday 1/8/02

Andy B arrives which is more than can be said for his luggage. Tonight's our official first night but we don't start until 2100 to pay back what we pinched the night before. We have a very good night. Weather is good, sseing is average (~1.1 arcsec fwhm in J) and the telescope works very well. We look at a few standard stars, Uranus, M30, the PN NGC6369, a ULIRG and a mysterious nebula near the galactic centre that we originally discovered with CIRSI. We work out the exact scales on the IFU. The 0.36 arcsec scale turns out be 0.36 arcsec and the 0.25 arcsec scale turns out to be 0.25 arcsec per lens - it does exactly what it says on the tin.

Friday 2/8/02

We look at more stuff but the weather's not so good and we close the dome at 0100.

Saturday 3/8/02

Gemini South Pictures

CIRPASS home page